# A Weighted Hankel Approach and Cramér-Rao Bound Analysis for Quantitative Acoustic Microscopy Imaging

### v1

@@ 3.2 ADMM-based solution
$\boldsymbol{\Lambda}^{q+1}\!$ $=\!\boldsymbol{\Lambda}^{q}+{\rho}(\boldsymbol{H}^{q+1}-H(\boldsymbol{g}^{q+1})),$ (11)
@@
$\mathcal{L}(\boldsymbol{H},\boldsymbol{g},$ $\boldsymbol{\Lambda})\!=\!\mathcal{R}_{P}(\boldsymbol{H})+\frac{1}{2}{\boldsymbol{w^{T}}}||\boldsymbol{x}-\boldsymbol{g}||_{2}^{2}$ (12)
$+\langle\boldsymbol{\Lambda},\boldsymbol{H}-H(\boldsymbol{g})\rangle_{\text{Re}}+\frac{\rho}{2}||\boldsymbol{H}-H(\boldsymbol{g})||_{F}^{2},$
@@
is the augmented Lagrangian associated with (8). In (12), $\rho$ is a penalty coefficient, $\langle\boldsymbol{A},\boldsymbol{B}\rangle_{\text{Re}}\!=\!\operatorname{real}(\operatorname{tr}(\boldsymbol{A}\boldsymbol{B}^{H}))$ denotes the inner product between the matrices $\boldsymbol{A}$ and $\boldsymbol{B}$, $\cdot^{H}$ is the conjugate transpose, and $||\boldsymbol{A}||^{2}_{F}={\langle\boldsymbol{A},\boldsymbol{A}\rangle}_{\text{Re}}$ is the Frobenius norm.
@@
The ADMM procedure defined in (9)-(10)-(11) ends after a finite number $Q$ of iterations. The rank constraint makes (8) a non-convex problem, meaning that convergence to a unique solution is not guaranteed and may depend on the choice of parameters $\rho$ and $Q$ as well as the initial values $\boldsymbol{g}^{0}$ and $\boldsymbol{\Lambda}^{0}$. Nevertheless, results in [18] show that the HK-based spectral method performs substantially better than established high-resolution techniques like ESPRIT [24]. In this work, we fixed the parameters to $Q\!=\!200$, $\rho\!=\!0.025$, $\boldsymbol{g}^{0}\!=\!\boldsymbol{x}$ and $\boldsymbol{\Lambda}^{0}\!=\!\boldsymbol{0}$, following the configuration in [18, 19].
@@
The most computationally intensive operation in each iteration step $q$ is the singular value decomposition (SVD) performed during the update of $\boldsymbol{H}$. The time complexity of the SVD operation is $\mathcal{O}(N^{3})$, leading to an overall time complexity for the algorithm of $\mathcal{O}(QN^{3})$. In comparison, the ESPRIT method requires the computation of only a single SVD, while the AR approach does not require any SVD computations.
@@
Due to the regularization term $\frac{\rho}{2}||\boldsymbol{H}-H(\boldsymbol{g})||_{F}^{2}$ in (12), it is common that $\boldsymbol{\hat{H}}\!\neq\!H(\boldsymbol{\hat{g}})$, where $\boldsymbol{\hat{H}}$ and $\boldsymbol{\hat{g}}$ are estimates of $\boldsymbol{H}$ and $\boldsymbol{g}$ resulting from the ADMM algorithm. This discrepancy suggests that $\boldsymbol{\hat{g}}$ or the vector $\boldsymbol{\hat{x}}^{\text{aprox}}$, obtained by averaging the anti-diagonals terms of $\boldsymbol{\hat{H}}$ provides a valid approximation of a sum of exponential functions and can serve as a solution to (8). The vector $\boldsymbol{\hat{x}}^{\text{aprox}}$ is defined element-wise as
$x^{\text{aprox}}_{\kappa_{n}}\!=\!\frac{1}{\mu_{n}}\underset{i+j=n+1}{\sum}[\boldsymbol{\hat{H}}]_{ij},$ (13)
@@
where $\mu_{n}$ is the number of times that $\boldsymbol{\hat{g}}_{\kappa_{n}}$ appears in the matrix $\boldsymbol{\hat{H}}$. Finally, the latter $\boldsymbol{x}^{\text{aprox}}$ is considered a more appropriate solution because $\boldsymbol{\hat{g}}$ is computed to best approximate the noisy data.
@@ 3.3 Spectral parameter estimation
By construction, the Hankel matrix solution $\boldsymbol{H}\!=\!H(\boldsymbol{x}^{\text{aprox}})$ is a complex symmetric matrix which satisfies $\boldsymbol{H}\!=\!\boldsymbol{H}^{T}$ and $\operatorname{rank}(\boldsymbol{H})\!=\!P$. For matrices with these properties, one can choose the matrices $\boldsymbol{U}$ and $\boldsymbol{V}$ appearing in the SVD of $\boldsymbol{H}$, i.e., $\boldsymbol{H}\!=\!\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{H}$ such that $\boldsymbol{U}\!=\!\overline{\boldsymbol{V}}$, where the bar indicates complex conjugation. Consequently,
$\boldsymbol{H}\!=\!\sum_{p=1}^{P}{s}_{p}\boldsymbol{u}_{p}\boldsymbol{u}_{p}^{T},\,s_{p}\in\mathbb{R}^{+},\,\boldsymbol{u}_{p}\in\mathbb{C}^{2N+1}$ (14)
where ${s}_{1},{s}_{2},\ldots,{s}_{P}>0$ are the eigenvalues of $\boldsymbol{H}$. Moreover, the con-eigenvectors $\boldsymbol{u}_{p}$ are orthogonal, satisfy the relation $\boldsymbol{H}\overline{\boldsymbol{u}}_{p}\!=\!s_{p}\boldsymbol{u}_{p}$ and can be expressed as the sum of the same $P$ exponentials of $\boldsymbol{x}^{\text{aprox}}$.
Hence, we can decompose the matrix $\boldsymbol{U}\!=\!(\boldsymbol{u}_{1},\ldots,\boldsymbol{u}_{P})$ as $\boldsymbol{U}\!=\!\boldsymbol{E}\boldsymbol{G}$, where $\boldsymbol{G}$ is some invertible $P\times P$ matrix and $\boldsymbol{E}$ is a $2N+1\times P$ Vandermond-like matrix defined by $[\boldsymbol{E}]_{np}=\lambda_{p}^{\kappa_{n}}$, for $p\!=\!1,\ldots,P$ and $n\!=\!1,\ldots,2N+1$, i.e.,
@@
3.4 Iteratively re-weighted approach
@@
The proposed iteratively reweighting strategy starts with initial weights set to 1 (no weighting). After performing spectral estimation, the residuals from the spectral reconstruction are calculated, and the weights are updated according to (17). This iterative process gradually down-weights outliers (data points with high residuals), thereby reducing their influence on the estimation and improving robustness in the presence of noise.
@@ 4 Cramér-Rao Bounds
The Cramér-Rao bound (CRB) [29] is a theoretical lower bound on the variance of any unbiased estimator for a given parametric model. The CRB provides a valuable benchmark for assessing the performance of estimation algorithms, as it defines the best achievable performance given the statistical characteristics of the data and the noise. In addition, deriving this bound offers insight into the estimation problem by identifying which parameters are more difficult to estimate and under what conditions. It also helps corroborate the underlying physical theory and establish experimental conditions that enable accurate parameter estimation.
@@ 4.1 Bounds for spectral parameter estimation
The CRB for the real-valued parameter vector $\boldsymbol{\theta}\!=\!(A_{1},\ldots,A_{P},b_{1},\ldots,b_{P},\gamma_{1},\ldots,\gamma_{P},\nu_{1},\ldots,\nu_{P})$ in spectral model (4) is defined by the diagonal elements of the inverse of the Fisher Information Matrix (FIM) $\boldsymbol{F}(\boldsymbol{\theta})$ [29], i.e.,
@@
Following the presentations in [30] and [31], we use the derivatives $\dfrac{\partial g_{\kappa}}{\partial{A_{p}}}\!=\!\exp{(\mathrm{i}b_{p})}\lambda_{p}^{\kappa}$, $\dfrac{\partial g_{\kappa}}{\partial{\beta_{p}}}\!=\!\mathrm{i}a_{p}\lambda_{p}^{\kappa}$, $\dfrac{\partial g_{\kappa}}{\partial{\gamma_{p}}}\!=\!2\pi\kappa_{p}\lambda_{p}^{\kappa}$ and $\dfrac{\partial g_{\kappa}}{\partial{\nu_{p}}}\!=\!\mathrm{i}2\pi\kappa a_{p}\lambda_{p}^{\kappa}$ to form the matrix
@@
and $\boldsymbol{B}\!=\!\operatorname{diag}\{(\exp{(\mathrm{i}b_{1})},\ldots,\exp{(\mathrm{i}b_{P})})\}$. Therefore, we can express the inverse FIM as
@@
where $\boldsymbol{P}=\operatorname{diag}[\boldsymbol{I}\,\boldsymbol{A}\,\boldsymbol{A}\,\boldsymbol{A}]$ with $\boldsymbol{A}\!=\!\operatorname{diag}\{(a_{1},\ldots,a_{p})\}$.
@@ 4.2 Bounds for acoustic parameter estimation
where $\nabla_{\theta}[\boldsymbol{\phi}^{T}(\boldsymbol{\theta})]=\left(\dfrac{\partial\boldsymbol{\phi}_{1}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}},\dfrac{\partial\boldsymbol{\phi}_{2}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}},\dfrac{\partial\boldsymbol{\phi}_{3}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}},\dfrac{\partial\boldsymbol{\phi}_{4}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\right)$ is the Jacobian matrix of $\boldsymbol{\phi}^{T}(\boldsymbol{\theta})$ with respect to $\boldsymbol{\theta}$. The partial derivatives required to construct the Jacobian are given by
@@
$\dfrac{\partial\boldsymbol{\phi}_{4}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}$ $=(2Z_{w}/(R_{wg}(1-a_{1}/R_{wg})^{2},\boldsymbol{0}_{4P-1}),$
@@ 5 Material and Methods
Current QAM systems are capable of operating at frequencies of 250 MHz [6] and 500 MHz [32]. Although increasing the transducer frequency improves spatial resolution, it also introduces heightened sensitivity to environmental factors such as vibrations and temperature fluctuations, increasing the challenges in the estimation of acoustic parameters.
The HK-based methods introduced in Section 3, referred to as RHK for the proposed reweighted approach and HK for the nonweighted variant [19] (which corresponds to the first iteration of RHK), along with the standard AR-based algorithm [15], are applicable across different transducer frequencies. In this study, we focus on real and simulated QAM data acquired using a 500 MHz transducer.
In QAM, estimates are considered outliers if they fall outside physically admissible ranges, specifically $c\!<\!1500$ m/s, $c\!>\!2200$ m/s, $Z\!<\!1.48$ MRayl, or $Z\!>\!2.2$ MRayl [15]. Therefore, beyond assessing estimator accuracy, the outlier count serves as an additional metric to evaluate the efficacy of each method.
@@ 5.1 Simulation setting
For a given set of acoustic parameters, we simulated QAM RF signals applying the inverse Fourier transform to the model in (3), setting $P\!=\!2$, $M\!=\!300$, and $f_{s}\!=\!10\,\text{GHz}$. The resulting signals were then perturbed by independent and identically distributed (i.i.d.) zero mean Gaussian noise with variance $\sigma_{x}^{2}$, given by
@@
where $\xi(h_{0})$ is the maximum of the absolute value of the envelope of $h_{0}$, estimated using its Hilbert transform-based analytic signal [15]. This noise formulation ensures that the simulated signals maintain the desired signal-to-noise ratio (SNR) in the time domain, representing the difference in decibels between the amplitude of the reference signal and the added noise. To assess estimation performance, we generated 200 realizations of the random Gaussian noise for each parameter configuration.
To reflect real-world conditions, we used a measured reference signal $h_{0}$ from our 500 MHz QAM system [32] and set the acoustic parameters to experimentally relevant values for QAM applications. These values were chosen based on preliminary tests to determine the optimal range between easily separable scenarios (e.g., large SNR, $d$ and $|Z-Z_{w}|$, and small $\alpha$) with more challenging cases (e.g., small SNR, $d$, and $|Z-Z_{w}|$, and large $\alpha$[15].
In each simulation, we varied the parameter under investigation while keeping the remaining parameters fixed as follows: SNR $\!=\!50$ dB, $d\!=\!4$ $\mu$m, $Z\!=\!1.63$ MRayl, $c\!=\!1600$ m/s, $\alpha\!=\!10$ dB/MHz/cm. We evaluated the effect of decreasing the SNR, reducing signal separation (i.e., thinner samples), lowering the amplitude of the first signal (i.e., acoustic impedance contrast $|Z-Z_{w}|$), and increasing sample attenuation. Table 1 summarizes the ranges of all parameter values used in our simulations.

### v2

@@ 3.2 ADMM-based solution
$\boldsymbol{\Lambda}^{q+1}\!$ $=\!\boldsymbol{\Lambda}^{q}+{\rho}(\boldsymbol{H}^{q+1}-\mathcal{H}(\boldsymbol{g}^{q+1})),$ (12)
@@
$\mathcal{L}(\boldsymbol{H},\boldsymbol{g},\boldsymbol{\Lambda})\!=$ $\!\mathcal{R}_{P}(\boldsymbol{H})+{\frac{1}{2}\,\|\boldsymbol{x}-\boldsymbol{g}\|_{\boldsymbol{W}}^{2}}$ (13)
$+\langle\boldsymbol{\Lambda},\boldsymbol{H}-\mathcal{H}(\boldsymbol{g})\rangle_{\text{Re}}+\frac{\rho}{2}||\boldsymbol{H}-\mathcal{H}(\boldsymbol{g})||_{F}^{2},$
@@
is the augmented Lagrangian associated with (9). In (13), $\rho$ is a penalty coefficient, $\langle\boldsymbol{A},\boldsymbol{B}\rangle_{\text{Re}}\!=\!\operatorname{real}(\operatorname{tr}(\boldsymbol{A}\boldsymbol{B}^{H}))$ denotes the inner product between the matrices $\boldsymbol{A}$ and $\boldsymbol{B}$, and $||\boldsymbol{A}||^{2}_{F}={\langle\boldsymbol{A},\boldsymbol{A}\rangle}_{\text{Re}}$ is the Frobenius norm.
@@
The ADMM procedure defined in (10)-(11)-(12) ends after a finite number $Q$ of iterations. The rank constraint makes (9) a non-convex problem, meaning that convergence to a unique solution is not guaranteed and may depend on the choice of parameters $\rho$ and $Q$ as well as the initial values $\boldsymbol{g}^{0}$ and $\boldsymbol{\Lambda}^{0}$. Nevertheless, results in [28] show that the Hankel-based spectral method performs substantially better than established high-resolution techniques like ESPRIT [37]. In this work, we fixed the parameters to $Q\!=\!200$, $\rho\!=\!0.025$, $\boldsymbol{g}^{0}\!=\!\boldsymbol{x}$ and $\boldsymbol{\Lambda}^{0}\!=\!\boldsymbol{0}$, following the configuration in [28, 38].
@@
The most computationally intensive operation in each iteration step $q$ is the singular value decomposition (SVD) performed during the update of $\boldsymbol{H}$. The time complexity of the SVD operation is $\mathcal{O}(N^{3})$, leading to an overall time complexity for the algorithm of $\mathcal{O}(QN^{3})$. In comparison, the ESPRIT method requires the computation of only a single SVD, while the AR approach [25] does not require any SVD computations.
@@
Due to the regularization term $\frac{\rho}{2}||\boldsymbol{H}-\mathcal{H}(\boldsymbol{g})||_{F}^{2}$ in (13), it is common that $\boldsymbol{\hat{H}}\!\neq\!\mathcal{H}(\boldsymbol{\hat{g}})$, where $\boldsymbol{\hat{H}}$ and $\boldsymbol{\hat{g}}$ are estimates of $\boldsymbol{H}$ and $\boldsymbol{g}$ resulting from the ADMM algorithm. This discrepancy suggests that $\boldsymbol{\hat{g}}$ or the vector ${\boldsymbol{x}}^{\text{approx}}$, obtained by averaging the anti-diagonals terms of $\boldsymbol{\hat{H}}$ provides a valid approximation of a sum of exponential functions and can serve as a solution to (9). The vector $\boldsymbol{x}^{\text{approx}}$ is defined element-wise as
$x^{\text{approx}}_{\kappa_{n}}\!=\!\frac{1}{\mu_{n}}\underset{i+j=n+1}{\sum}[\boldsymbol{\hat{H}}]_{ij},$ (14)
@@
where $\mu_{n}$ is the number of times that ${\hat{g}}_{\kappa_{n}}$ appears in the matrix $\boldsymbol{\hat{H}}$. Finally, the latter $\boldsymbol{x}^{\text{approx}}$ is considered a more appropriate solution because $\boldsymbol{\hat{g}}$ is computed to best approximate the noisy data.
@@ 3.3 Spectral parameter estimation
By construction, the Hankel matrix solution $\boldsymbol{H}\!=\!\mathcal{H}(\boldsymbol{x}^{\text{approx}})$ is a complex symmetric matrix which satisfies $\boldsymbol{H}\!=\!\boldsymbol{H}^{T}$ and $\operatorname{rank}(\boldsymbol{H})\!=\!P$. For matrices with these properties, one can choose the matrices $\boldsymbol{U}$ and $\boldsymbol{V}$ appearing in the SVD of $\boldsymbol{H}$, i.e., $\boldsymbol{H}\!=\!\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{H}$ such that $\boldsymbol{U}\!=\!\overline{\boldsymbol{V}}$, where the bar indicates complex conjugation. Consequently,
$\boldsymbol{H}\!=\!\sum_{p=1}^{P}{s}_{p}\boldsymbol{u}_{p}\boldsymbol{u}_{p}^{T},\,s_{p}\in\mathbb{R}^{+},\,\boldsymbol{u}_{p}\in\mathbb{C}^{2N+1},$ (15)
where ${s}_{1},{s}_{2},\ldots,{s}_{P}>0$ are the eigenvalues of $\boldsymbol{H}$. Moreover, the con-eigenvectors $\boldsymbol{u}_{p}$ are orthogonal, satisfy the relation $\boldsymbol{H}\overline{\boldsymbol{u}}_{p}\!=\!s_{p}\boldsymbol{u}_{p}$ and can be expressed as the sum of the same $P$ exponentials of $\boldsymbol{x}^{\text{approx}}$.
Hence, we can decompose the matrix $\boldsymbol{U}\!=\!(\boldsymbol{u}_{1},\ldots,\boldsymbol{u}_{P})$ as $\boldsymbol{U}\!=\!\boldsymbol{E}\boldsymbol{G}$, where $\boldsymbol{G}$ is some invertible $P\times P$ matrix and $\boldsymbol{E}$ is a $2N+1\times P$ Vandermonde-like matrix defined by $[\boldsymbol{E}]_{np}=\lambda_{p}^{\kappa_{n}}$, for $p\!=\!1,\ldots,P$ and $n\!=\!1,\ldots,2N+1$, i.e.,
@@
3.4 Iteratively reweighted approach
@@
The proposed iteratively reweighting strategy starts with initial weights set to 1. After performing spectral estimation, the residuals from the spectral reconstruction are calculated, and the weights are updated according to (18). This iterative process gradually down-weights data points with high residuals, thereby reducing their influence on the estimation and improving robustness in the presence of noise.
@@ 4 Cramér-Rao Bounds
The Cramér-Rao bound (CRB) [50] is a theoretical lower bound on the variance of any unbiased estimator for a given parametric model. The CRB provides a valuable benchmark for assessing the performance of estimation algorithms, as it defines the best achievable performance given the statistical characteristics of the data and the noise. In addition, deriving this bound offers insight into the estimation problem by identifying which parameters are more difficult to estimate and under what conditions. It also helps corroborate the underlying physical hypothesis and establish experimental conditions that enable accurate parameter estimation.
@@ 4.1 Bounds for spectral parameter estimation
The CRB for the real-valued parameter vector
$\boldsymbol{\theta}\!=\!(A_{1},\ldots,A_{P},b_{1},\ldots,b_{P},\gamma_{1},\ldots,\gamma_{P},\nu_{1},\ldots,\nu_{P})$
in spectral model (5) is defined by the diagonal elements of the inverse of the Fisher Information Matrix (FIM) $\boldsymbol{F}(\boldsymbol{\theta})$ [50], i.e.,
@@
Following the presentations in [51] and [52], we use the derivatives
$\dfrac{\partial g_{\kappa}}{\partial{A_{p}}}\!=\!\exp{(\mathrm{i}b_{p})}\lambda_{p}^{\kappa},\,\dfrac{\partial g_{\kappa}}{\partial{\beta_{p}}}\!=\!\mathrm{i}a_{p}\lambda_{p}^{\kappa},\,\dfrac{\partial g_{\kappa}}{\partial{\gamma_{p}}}\!=\!2\pi\kappa_{p}\lambda_{p}^{\kappa}\text{\, and \,}\dfrac{\partial g_{\kappa}}{\partial{\nu_{p}}}\!=\!\mathrm{i}2\pi\kappa a_{p}\lambda_{p}^{\kappa}$
to form the matrix
@@
and $\boldsymbol{B}\!=\!\operatorname{diag}(\exp{(\mathrm{i}b_{1})},\ldots,\exp{(\mathrm{i}b_{P})})$. Therefore, we can express the inverse FIM as
@@
where $\boldsymbol{P}=\operatorname{diag}[\boldsymbol{I}\,\boldsymbol{A}\,\boldsymbol{A}\,\boldsymbol{A}]$ with $\boldsymbol{A}\!=\!\operatorname{diag}(a_{1},\ldots,a_{p})$.
@@ 4.2 Bounds for acoustic parameter estimation
where $\nabla_{\theta}[\boldsymbol{\phi}^{T}(\boldsymbol{\theta})]=\left(\dfrac{\partial\boldsymbol{\phi}_{1}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}},\dfrac{\partial\boldsymbol{\phi}_{2}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}},\dfrac{\partial\boldsymbol{\phi}_{3}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}},\dfrac{\partial\boldsymbol{\phi}_{4}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\right)$ is the Jacobian matrix of $\boldsymbol{\phi}^{T}(\boldsymbol{\theta})$ with respect to $\boldsymbol{\theta}$, and
@@
$\dfrac{\partial\boldsymbol{\phi}_{4}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}$ $=(2Z_{w}/(R_{wg}(1-a_{1}/R_{wg}))^{2},\boldsymbol{0}_{4P-1}),$
@@ 5 Material and Methods
QAM systems are capable of operating at frequencies of 250 MHz [8] and 500 MHz [53]. Although increasing the transducer frequency improves spatial resolution, it also introduces heightened sensitivity to environmental factors such as vibrations and temperature fluctuations, increasing the challenges in the estimation of acoustic parameters.
The Hankel-based methods introduced in Section 3, referred to as RHK for the proposed reweighted approach and HK for the non-weighted version [38] (which corresponds to the initialization of RHK), along with the standard autoregressive algorithm [25], are applicable across different transducer frequencies. In this study, we focus on real and simulated QAM data acquired using a 500 MHz transducer.
QAM estimates are considered outliers or unreliable values if they fall outside physically admissible ranges, specifically $c\!<\!1500$ m/s, $c\!>\!2200$ m/s, $Z\!<\!1.48$ MRayl, or $Z\!>\!2.2$ MRayl [25]. Therefore, beyond assessing estimator accuracy using the root mean square error (RMSE), the outlier count serves as an additional metric to evaluate the efficacy of each method. From this point onward, outliers are excluded from RMSE calculations to ensure a fair comparison across methods.
@@ 5.1 Simulation setting
We used a measured reference signal $h_{0}$ from our 500 MHz QAM system, to simulate QAM RF signals applying the inverse Fourier transform to the model in (4), for a given set of acoustic parameters, $P\!=\!2$, $M\!=\!300$, and $f_{s}\!=\!10\,\text{GHz}$. The resulting signals were then perturbed by independent and identically distributed (i.i.d.) zero mean Gaussian noise with variance $\sigma_{x}^{2}$, given by
@@
where $\xi(h_{0})$ is the maximum of the absolute value of the envelope of $h_{0}$, estimated using its Hilbert transform-based analytic signal [25]. This noise formulation ensures that the simulated signals maintain the desired signal-to-noise ratio (SNR) in the time domain, representing the difference in decibels (dB) between the amplitude of the reference signal and the added noise. To assess estimation performance, we generated 200 realizations of the random Gaussian noise for each parameter configuration.
In each simulation, the acoustic parameters were set to experimentally relevant values for QAM applications, chosen based on preliminary tests to determine the optimal range between easily separable scenarios (e.g., large SNR, $d$ and $|Z-Z_{w}|$, and small $\alpha$) with more challenging cases (e.g., small SNR, $d$, and $|Z-Z_{w}|$, and large $\alpha$[25]. We varied the parameter under investigation while keeping the remaining parameters fixed as follows: SNR $\!=\!50$ dB, $d\!=\!4$ $\mu$m, $Z\!=\!1.63$ MRayl, $c\!=\!1600$ m/s, $\alpha\!=\!10$ dB/MHz/cm. We evaluated the effect of decreasing the SNR, reducing signal separation (i.e., thinner samples), lowering the amplitude of the first signal (i.e., acoustic impedance contrast $|Z-Z_{w}|$), and increasing sample attenuation. Table 1 summarizes the ranges of all parameter values used in our simulations.